# Safely Learning Controlled Stochastic Dynamics

### v2

@@ 1 Introduction
We consider the problem of safely learning the dynamics of controlled continuous-time stochastic systems from discrete-time observations of trajectory data. This setting is common in applications such as robotics, finance, and healthcare, where system dynamics are only partially known and must be estimated from data. A key challenge in these applications is ensuring safety during both the learning phase and subsequent deployment [bonalli2022; lew2024]. As an example, consider an autonomous robot navigating a partially known and turbulent environment, as illustrated in Figure 1. While the deterministic part of the dynamics may be approximately modeled using prior knowledge, the stochastic disturbances (represented by the brown region in Figure 1) due to wind or sensor noise are often unknown and must be learned. Collecting data through naive exploration can result in unsafe trajectories, potentially causing damage to the system or its environment. A second example arises in financial portfolio management, where the drift component of asset prices may be known from historical data, but market volatility remains uncertain. Safety here may correspond to the requirement that the portfolio value stays above a critical threshold with high probability, simulating portfolio loss aversion in risk-sensitive financial decision-making. These examples highlight a common need: learning stochastic dynamics of a system from data, while ensuring safety throughout the process. This requires ensuring that all executed trajectories remain within a predefined safe region with high probability. In addition, at deployment time, the learned model should enable prediction of whether a proposed control input satisfies the safety requirements, including those not encountered during training [wabersich2021predictive; lindemann2023safe].
@@
  • Provably safe exploration and adaptive estimation rates. We prove that the proposed method guarantees safety and derive learning rates for system estimation, with rates that are adaptive to the Sobolev regularity of the underlying dynamics. Crucially, our approach requires only smooth dynamics (with respect to time, state, and control variables) and an initial non-empty set of known safe controls. These mild assumptions make the method applicable to a broad range of complex real-world systems subject to stochastic disturbances, found in diverse areas including robotics, fluid flow control, and chemical reaction control [martin2010true; plappert2018multi; morton2018deep; peters2022causal; fu2024mobile; wei2024diffphycon].
@@ 2 Background
Let $X$ be a dynamical system governed by a non-linear, $n$-dimensional controlled SDE [bonalli2023; lew2023],
@@
captures second-order dynamics under control $u$ and state-dependent diffusion $a\bigl(X(t)\bigr)$. Real-world examples include drones navigating turbulent environments (where $u$ represents thrust or steering, and $a(\cdot)$ models wind turbulence), fluid-dynamical systems [batchelor2000introduction], and molecular dynamics (where $u$ captures external forces such as optical traps, and $a(\cdot)$ reflects spatially varying thermal fluctuations [volpe2013simulation; marago2013optical; pesce2020optical]).
@@
Let $X_{u}$ denote the solution to Eq. (1) under the control $u\in\mathcal{H}$ (well-defined by existence and uniqueness; see Sec.3 in bonalli2025non). Let $u_{\theta}:[0,T_{\max}]\times\mathbb{R}^{n}\to\mathbb{R}^{d}$ denote a family of controls parameterized by $\theta\in D\subset\mathbb{R}^{m}$, where $D$ is a compact subset of $\mathbb{R}^{m}$. Such a finite-dimensional parameterization is a mild assumption that is widely prevalent in many real-world applications, including robotics [berkenkamp2017safe], process control [seborg2016process], and financial engineering [merton1975optimum].
@@
where $p_{\theta}(t,x)$ is the density of the state $X_{u_{\theta}}(t)$ under control $u_{\theta}$, well-defined by standard existence and uniqueness results (see Sec.3 in bonalli2025non). To estimate this density, we collect a dataset of trajectories
@@ 2.1 Related work
Safe learning in control systems under uncertainty is a central topic in reinforcement learning, with methods built on assumptions such as known dynamics [hewing2020learning], controllability [turchetta2016safe; mania2022active; selim2022safe], or recovery policies [hans2008safe; moldovan2012safe].
Much of the literature focuses on discrete-time or discrete-state systems modeled as Markov Decision Processes (MDPs), where risk-sensitive and safe exploration techniques have been developed [coraluppi1999risk; geibel2005risk; moldovan2012safe; garcia2012safe]. Among these, Safe Bayesian Optimization (BO) methods stand out for providing some of the strongest high-confidence safety guarantees during exploration [sui2015safe; sui2018stagewise; bottero2022information; li2024safe], particularly in MDP settings with known dynamics or access to safety level evaluations [turchetta2016safe; turchetta2019safe]. In continuous domains, early work focused on designing control policies that avoid unsafe regions [akametalu2014reachability; koller2018learning], while more recent approaches incorporate offline learning and online adaptation for nonlinear systems [lew2022safe]. Stability-based guarantees via Lyapunov theory offer formal certification but often require full dynamics knowledge [berkenkamp2017safe; richards2018lyapunov]. Safe BO has also been adapted to continuous settings [sukhija2023gosafeopt; prajapat2024safe], under assumptions such as access to dynamics, safety oracles, or specific control-theoretic properties. Joint estimation of both dynamics and safety remains comparatively underexplored, particularly in continuous-time settings [ahmadi2021safely; dalal2018safe; wabersich2018safe; fisac2018general; wabersich2018linear; khojasteh2020probabilistic; wagener2021safe; wachi2020safe; gu2022review].
@@ 3 Assumptions
As formalized by the No-Free-Lunch Theorem [devroye2013probabilistic], learning is only possible under prior assumptions. We now state and discuss the key assumptions used in this work.
@@
This assumption ensures that at least one control is known to be safe at the outset, allowing safe exploration to begin. In fact, $S_{0}$ may be as small as a singleton; only one known safe control is required. Without such a point, safe learning cannot be initiated. We express the assumption in set form to allow for larger safe sets, which can accelerate exploration while preserving guarantees. This is a standard assumption in the literature of safe UCB methods [sui2015safe; turchetta2016safe; berkenkamp2017safe], and is realistic in many applications including robotics [sukhija2023gosafeopt] and safety-critical process control [qin2003survey], where systems naturally start in safe conditions, e.g., $S_{0}=\{(0,\theta):\theta\in D\}$.
@@
This assumption guarantees the existence of at least one control capable of returning the system to the reset region with high probability. Assumption (A2) enables the generation of independent sample paths starting from the same initial conditions, which is crucial for evaluating variance and managing uncertainty during safe exploration. As with Assumption (A1), one known reset point is sufficient, though larger reset sets accelerate learning. A simple case is $R_{0}=\{(0,\theta):\theta\in D\}$, corresponding to systems that can always be reset from the initial condition. In practice, reset feasibility depends on system constraints: for instance, batch chemical reactors can often be reset only in early phases, before irreversible reactions occur [seborg2016process]. In contrast, many autonomous systems like drones or driving robots can usually be reset, at least during training.
@@
This smoothness assumption ensures that the system dynamics are sufficiently regular for our purposes. It is standard in statistical learning theory and underpins our convergence guarantees [pillaud2018statistical]. Sobolev regularity of the drift and diffusion terms in the underlying stochastic differential equation is expected to imply Sobolev regularity of the resulting state densities under standard conditions. This follows from classical results in parabolic PDE theory, where solutions typically gain regularity relative to the coefficients, roughly two derivatives in space and one in time. A formal analysis of this connection is beyond the scope of the present work and is left for future investigation (see bonalli2025non for related results).
@@ 4 Proposed method
We propose a method for safely exploring and learning system dynamics over a parameterized control space $\mathcal{H}=\{u_{\theta}\mid\theta\in D\subset\mathbb{R}^{m}\}$. Following the safe UCB framework [sui2015safe; sui2018stagewise; bottero2022information], our goal is to select controls that reduce model uncertainty while ensuring, with high probability, that trajectories (i) remain within the safe region and (ii) end in the reset region. We jointly learn three models: a dynamics model (state densities), a safety model (safety probabilities), and a reset model (reset probabilities), each equipped with confidence bounds from a shared kernel. This enables active exploration under high-probability constraints. After training, the learned models support inference on unseen inputs and yield a certified control set that can be deployed with safety guarantees.
@@ 4.2 System estimation
where $\rho_{R}(x)\triangleq R^{n/2}\|x\|^{-n/2}B_{n/2}(2\pi R\|x\|)$, $R>0$, and $B_{n/2}$ is the Bessel $J$ function of order $n/2$ (See bonalli2025non).
@@ 4.3 Safe sampling
The derivation in Appendix A.7 shows that the confidence parameters depend on upper bounds of the RKHS norms of the safety and reset functions. These bounds may be available from prior knowledge of the system’s regularity. We view this prior knowledge as a reasonable minimal assumption for guaranteeing safe learning under unknown dynamics. When unavailable, the bounds can be conservatively overestimated, ensuring safety but potentially leading to slower exploration. Developing adaptive strategies to estimate these quantities without prior knowledge is a promising direction for future work, for instance through online adaptation via the doubling trick [shalev2012online].
@@ 5 Safety and estimation guarantees for Sobolev dynamics
Safety and exploration guarantees for safe kernelized UCB methods have been developed in prior work [sui2015safe; sui2018stagewise; bottero2022information], grounded in kernelized bandit theory [srinivas2009gaussian; valko2013finite; janz2020bandit], which in turn builds on linear bandit results [dani2008stochastic; auer2002using]. Building on this foundation, we establish novel theoretical guarantees for safe exploration and dynamics estimation under Sobolev regularity. Complete proofs are deferred to Appendix A.
@@ A.3 Proof of the validity of confidence intervals
This assumption ensures that the safety and reset level functions can be represented in a suitable function space for estimation. In particular, it encodes prior knowledge; for example, if the functions are known to lie in a Sobolev (Hilbert) space $H^{m}$, one may choose a Sobolev kernel of order $m$. This is a standard assumption in the literature on kernel methods. In many practical applications, such as industrial process control and robotics [sukhija2023gosafeopt; qin2003survey], SDEs with smooth coefficients produce smooth probability densities [bonalli2025non], ensuring that $s(\theta,t)$ is smooth and can be accurately represented by Gaussian or Sobolev kernels. Therefore, this assumption is mild in practice.
@@
where $\beta_{N}^{s}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|+\|s\|_{\mathcal{G}}$, $\beta_{N}^{r}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{r}_{\theta_{i},t_{i}}-r(\theta_{i},t_{i})|+\|r\|_{\mathcal{G}}$.
@@
$\leq N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|\,\|\hat{C}^{1/2}\hat{C}_{\lambda}^{-1}\phi(\theta,t)\|_{\mathcal{G}}$ (39)
$\leq N^{-1/2}\lambda^{-1}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|\,\sigma_{N}(\theta,t).$ (40)
From Assumption (A4), $s(\theta,t)=\langle s,\phi(\theta,t)\rangle_{\mathcal{G}}$, such that $S=\Phi^{*}s$, and then $s_{N}(\theta,t)=s^{*}\Phi\Phi^{*}(\hat{C}+\lambda)^{-1}\phi(\theta,t)=s^{*}\hat{C}(\hat{C}+\lambda)^{-1}\phi(\theta,t)$.
@@
by defining $\beta_{N}^{p}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}\|\hat{p}_{\theta_{i},t_{i}}(\cdot)-p_{\theta_{i}}(t_{i},\cdot)\|_{L^{\infty}(\mathbb{R}^{n})}+\|p\|_{L^{\infty}(\mathbb{R}^{n};\mathcal{G})}$
@@
$(A)$ $\leq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{p}_{\theta_{i},t_{i}}(x)-p_{\theta_{i}}(t_{i},x)|\,\sigma_{N}(\theta,t).$ (50)
@@
by defining $\beta_{N}^{p}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}\|\hat{p}_{\theta_{i},t_{i}}(\cdot)-p_{\theta_{i}}(t_{i},\cdot)\|_{L^{\infty}(\mathbb{R}^{n})}+\|p\|_{L^{\infty}(\mathbb{R}^{n};\mathcal{G})}$

### v3

@@ 1 Introduction
We consider the problem of safely learning the dynamics of controlled continuous-time stochastic systems from discrete-time observations of trajectory data. This setting is common in applications such as robotics, finance, and healthcare, where system dynamics are only partially known and must be estimated from data. A key challenge in these applications is ensuring safety during both the learning phase and subsequent deployment [Bonalli et al., 2022; Lew et al., 2024]. As an example, consider an autonomous robot navigating a partially known and turbulent environment, as illustrated in Figure 1. While the deterministic part of the dynamics may be approximately modeled using prior knowledge, the stochastic disturbances (represented by the brown region in Figure 1) due to wind or sensor noise are often unknown and must be learned. Collecting data through naive exploration can result in unsafe trajectories, potentially causing damage to the system or its environment. A second example arises in financial portfolio management, where the drift component of asset prices may be known from historical data, but market volatility remains uncertain. Safety here may correspond to the requirement that the portfolio value stays above a critical threshold with high probability, simulating portfolio loss aversion in risk-sensitive financial decision-making. These examples highlight a common need: learning stochastic dynamics of a system from data, while ensuring safety throughout the process. This requires ensuring that all executed trajectories remain within a predefined safe region with high probability. In addition, at deployment time, the learned model should enable prediction of whether a proposed control input satisfies the safety requirements, including those not encountered during training [Wabersich and Zeilinger, 2021; Lindemann et al., 2023].
@@
  • Provably safe exploration and adaptive estimation rates. We prove that the proposed method guarantees safety and derive learning rates for system estimation, with rates that are adaptive to the Sobolev regularity of the underlying dynamics. Crucially, our approach requires only smooth dynamics (with respect to time, state, and control variables) and an initial non-empty set of known safe controls. These mild assumptions make the method applicable to a broad range of complex real-world systems subject to stochastic disturbances, found in diverse areas including robotics, fluid flow control, and chemical reaction control [Martin and Salaün, 2010; Plappert et al., 2018; Morton et al., 2018; Peters et al., 2022; Fu et al., 2024; Wei et al., 2024].
@@ 2 Background
Let $X$ be a dynamical system governed by a non-linear, $n$-dimensional controlled SDE [Bonalli and Bonnet, 2023; Lew et al., 2023],
@@
captures second-order dynamics under control $u$ and state-dependent diffusion $a\bigl(X(t)\bigr)$. Real-world examples include drones navigating turbulent environments (where $u$ represents thrust or steering, and $a(\cdot)$ models wind turbulence), fluid-dynamical systems [Batchelor, 2000], and molecular dynamics (where $u$ captures external forces such as optical traps, and $a(\cdot)$ reflects spatially varying thermal fluctuations [Volpe and Volpe, 2013; Marago et al., 2013; Pesce et al., 2020]).
@@
Let $X_{u}$ denote the solution to Eq. (1) under the control $u\in\mathcal{H}$ (well-defined by existence and uniqueness; see Sec.3 in Bonalli and Rudi [2025]). Let $u_{\theta}:[0,T_{\max}]\times\mathbb{R}^{n}\to\mathbb{R}^{d}$ denote a family of controls parameterized by $\theta\in D\subset\mathbb{R}^{m}$, where $D$ is a compact subset of $\mathbb{R}^{m}$. Such a finite-dimensional parameterization is a mild assumption that is widely prevalent in many real-world applications, including robotics [Berkenkamp et al., 2017], process control [Seborg et al., 2016], and financial engineering [Merton, 1975].
@@
where $p_{\theta}(t,x)$ is the density of the state $X_{u_{\theta}}(t)$ under control $u_{\theta}$, well-defined by standard existence and uniqueness results (see Sec.3 in Bonalli and Rudi [2025]). To estimate this density, we collect a dataset of trajectories
@@ 2.1 Related work
Safe learning in control systems under uncertainty is a central topic in reinforcement learning, with methods built on assumptions such as known dynamics [Hewing et al., 2020], controllability [Turchetta et al., 2016; Mania et al., 2022; Selim et al., 2022], or recovery policies [Hans et al., 2008; Moldovan and Abbeel, 2012].
Much of the literature focuses on discrete-time or discrete-state systems modeled as Markov Decision Processes (MDPs), where risk-sensitive and safe exploration techniques have been developed [Coraluppi and Marcus, 1999; Geibel and Wysotzki, 2005; Moldovan and Abbeel, 2012; Garcia and Fernández, 2012]. Among these, Safe Bayesian Optimization (BO) methods stand out for providing some of the strongest high-confidence safety guarantees during exploration [Sui et al., 2015, 2018; Bottero et al., 2022; Li et al., 2024], particularly in MDP settings with known dynamics or access to safety level evaluations [Turchetta et al., 2016, 2019]. In continuous domains, early work focused on designing control policies that avoid unsafe regions [Akametalu et al., 2014; Koller et al., 2018], while more recent approaches incorporate offline learning and online adaptation for nonlinear systems [Lew et al., 2022]. Stability-based guarantees via Lyapunov theory offer formal certification but often require full dynamics knowledge [Berkenkamp et al., 2017; Richards et al., 2018]. Safe BO has also been adapted to continuous settings [Sukhija et al., 2023; Prajapat et al., 2024], under assumptions such as access to dynamics, safety oracles, or specific control-theoretic properties. Joint estimation of both dynamics and safety remains comparatively underexplored, particularly in continuous-time settings [Ahmadi et al., 2021; Dalal et al., 2018; Wabersich and Zeilinger, 2018b; Fisac et al., 2018; Wabersich and Zeilinger, 2018a; Khojasteh et al., 2020; Wagener et al., 2021; Wachi and Sui, 2020; Gu et al., 2022].
@@ 3 Assumptions
As formalized by the No-Free-Lunch Theorem [Devroye et al., 2013], learning is only possible under prior assumptions. We now state and discuss the key assumptions used in this work.
@@
This assumption ensures that at least one control is known to be safe at the outset, allowing safe exploration to begin. In fact, $S_{0}$ may be as small as a singleton; only one known safe control is required. Without such a point, safe learning cannot be initiated. We express the assumption in set form to allow for larger safe sets, which can accelerate exploration while preserving guarantees. This is a standard assumption in the literature of safe UCB methods [Sui et al., 2015; Turchetta et al., 2016; Berkenkamp et al., 2017], and is realistic in many applications including robotics [Sukhija et al., 2023] and safety-critical process control [Qin and Badgwell, 2003], where systems naturally start in safe conditions, e.g., $S_{0}=\{(0,\theta):\theta\in D\}$.
@@
This assumption guarantees the existence of at least one control capable of returning the system to the reset region with high probability. Assumption (A2) enables the generation of independent sample paths starting from the same initial conditions, which is crucial for evaluating variance and managing uncertainty during safe exploration. As with Assumption (A1), one known reset point is sufficient, though larger reset sets accelerate learning. A simple case is $R_{0}=\{(0,\theta):\theta\in D\}$, corresponding to systems that can always be reset from the initial condition. In practice, reset feasibility depends on system constraints: for instance, batch chemical reactors can often be reset only in early phases, before irreversible reactions occur [Seborg et al., 2016]. In contrast, many autonomous systems like drones or driving robots can usually be reset, at least during training.
@@
This smoothness assumption ensures that the system dynamics are sufficiently regular for our purposes. It is standard in statistical learning theory and underpins our convergence guarantees [Pillaud-Vivien et al., 2018]. Sobolev regularity of the drift and diffusion terms in the underlying stochastic differential equation is expected to imply Sobolev regularity of the resulting state densities under standard conditions. This follows from classical results in parabolic PDE theory, where solutions typically gain regularity relative to the coefficients, roughly two derivatives in space and one in time. A formal analysis of this connection is beyond the scope of the present work and is left for future investigation (see Bonalli and Rudi [2025] for related results).
@@ 4 Proposed method
We propose a method for safely exploring and learning system dynamics over a parameterized control space $\mathcal{H}=\{u_{\theta}\mid\theta\in D\subset\mathbb{R}^{m}\}$. Following the safe UCB framework [Sui et al., 2015, 2018; Bottero et al., 2022], our goal is to select controls that reduce model uncertainty while ensuring, with high probability, that trajectories (i) remain within the safe region and (ii) end in the reset region. We jointly learn three models: a dynamics model (state densities), a safety model (safety probabilities), and a reset model (reset probabilities), each equipped with confidence bounds from a shared kernel. This enables active exploration under high-probability constraints. After training, the learned models support inference on unseen inputs and yield a certified control set that can be deployed with safety guarantees.
@@ 4.2 System estimation
where $\rho_{R}(x)\triangleq R^{n/2}\|x\|^{-n/2}B_{n/2}(2\pi R\|x\|)$, $R>0$, and $B_{n/2}$ is the Bessel $J$ function of order $n/2$ (See Bonalli and Rudi [2025]).
@@ 4.3 Safe sampling
The derivation in Appendix A.7 shows that the confidence parameters depend on upper bounds of the RKHS norms of the safety and reset functions. These bounds may be available from prior knowledge of the system’s regularity. We view this prior knowledge as a reasonable minimal assumption for guaranteeing safe learning under unknown dynamics. When unavailable, the bounds can be conservatively overestimated, ensuring safety but potentially leading to slower exploration. Developing adaptive strategies to estimate these quantities without prior knowledge is a promising direction for future work, for instance through online adaptation via the doubling trick [Shalev-Shwartz and others, 2012].
@@ 5 Safety and estimation guarantees for Sobolev dynamics
Safety and exploration guarantees for safe kernelized UCB methods have been developed in prior work [Sui et al., 2015, 2018; Bottero et al., 2022], grounded in kernelized bandit theory [Srinivas et al., 2009; Valko et al., 2013; Janz et al., 2020], which in turn builds on linear bandit results [Dani et al., 2008; Auer, 2002]. Building on this foundation, we establish novel theoretical guarantees for safe exploration and dynamics estimation under Sobolev regularity. Complete proofs are deferred to Appendix A.
@@ A.3 Proof of the validity of confidence intervals
This assumption ensures that the safety and reset level functions can be represented in a suitable function space for estimation. In particular, it encodes prior knowledge; for example, if the functions are known to lie in a Sobolev (Hilbert) space $H^{m}$, one may choose a Sobolev kernel of order $m$. This is a standard assumption in the literature on kernel methods. In many practical applications, such as industrial process control and robotics [Sukhija et al., 2023; Qin and Badgwell, 2003], SDEs with smooth coefficients produce smooth probability densities [Bonalli and Rudi, 2025], ensuring that $s(\theta,t)$ is smooth and can be accurately represented by Gaussian or Sobolev kernels. Therefore, this assumption is mild in practice.
@@
where $\beta_{N}^{s}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|+\|s\|_{\mathcal{G}}$, $\beta_{N}^{r}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{r}_{\theta_{i},t_{i}}-r(\theta_{i},t_{i})|+\|r\|_{\mathcal{G}}$.
@@
$\leq\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|\,\|\hat{C}^{1/2}\hat{C}_{\lambda}^{-1}\phi(\theta,t)\|_{\mathcal{G}}$ (39)
$\leq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|\,\sigma_{N}(\theta,t).$ (40)
From Assumption (A4), $s(\theta,t)=\langle s,\phi(\theta,t)\rangle_{\mathcal{G}}$, such that $S=\Phi^{*}s$, and then $s_{N}(\theta,t)=s^{*}\Phi\Phi^{*}(\hat{C}+\lambda I)^{-1}\phi(\theta,t)=s^{*}\hat{C}(\hat{C}+\lambda I)^{-1}\phi(\theta,t)$.
@@
by defining $\beta_{N}^{p}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}\|\hat{p}_{\theta_{i},t_{i}}(\cdot)-p_{\theta_{i}}(t_{i},\cdot)\|_{L^{\infty}(\mathbb{R}^{n})}+\|p\|_{L^{\infty}(\mathbb{R}^{n};\mathcal{G})}$
@@
$(A)$ $\leq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{p}_{\theta_{i},t_{i}}(x)-p_{\theta_{i}}(t_{i},x)|\,\sigma_{N}(\theta,t).$ (50)
@@
by defining $\beta_{N}^{p}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}\|\hat{p}_{\theta_{i},t_{i}}(\cdot)-p_{\theta_{i}}(t_{i},\cdot)\|_{L^{\infty}(\mathbb{R}^{n})}+\|p\|_{L^{\infty}(\mathbb{R}^{n};\mathcal{G})}$